OpenStudy (anonymous):

Suppose that y varies inversely with x. Write an equation for the inverse variation. y=2 when x=5

OpenStudy (anonymous):

Inverse variations are of the form y = k/x, where k is a constant. Plug your given values for x and y into this equation to find k: 2=k/5 ==> k=10. Thus the equation for the inverse variation is y=10/x